Circular No. 8923

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 E1 (CATALINA) An apparently asteroidal object discovery by the Catalina Sky Survey

(discovery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage, has been noted by numerous astrometric CCD observers to show cometary appearance. P. Birtwhistle (Great Shefford, Berkshire, England 0.40-m f/6 Schmidt-Cassegrain reflector, Mar. 3.82–3.83 UT, relatively poor seeing) writes that this object appears to be distinctly more diffuse than nearby stars of similar brightness, having a diameter of 9′′ and a possible extension from the center of light extending 12′′ in p.a. 40^{o}. J. V. Scotti (Spacewatch 1.8-m f/2.7 reflector, Mar. 4.12–4.16) reports a coma diameter of 6′′–8′′ and a tail extending ∼ 5′′.4–7′′.2 in p.a. 31^{o}. A. Boattini (Mt. Lemmon 1.5-m reflector, co-addition of four 90-s images from Mar. 4.13–4.14, bad seeing) measures a coma diameter of ∼ 7′′ × 8′′ (slight asymmetry along a NNE-SSW axis). E. Guido and G. Sostero (Castellammare di Stabia, Italy, remotely using a 0.25-m f/3.4 reflector near Mayhill, NM, U.S.A.; sixty co-added 60-s unfiltered exposures, Mar. 4.1) find a diffuse coma of diameter nearly 12′′.

2008 UT  α_{2000}  δ_{2000}  Mag.
Mar. 2.14405  4^{h}28^{m}12.^{s}21  −11^{o}26′17′′.6  19.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-E41.

T = 2008 Oct. 29.722 TT  ω = 285.^{o}589
Ω = 183.644 }2000.0
q = 4.82300 AU  i = 34.263

COMET C/2007 Y10 (SOHO) Further to IAUC 8921, tabulated below is the “discovery” observation

for an additional Kreutz sungrazing comet found on SOHO website images; the object appeared stellar (mag ∼ 5.5) in C3-coronagraph images, and showed a short, faint, diffuse tail in C2 images.

Comet  2007 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2007 Y10 Dec. 28.988 18^{h}41.^{m}5 −28^{o}06′ C3/2 BZ 2008-B61
2008 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green